# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ Abstract
Rhombohedral graphene multilayers host various broken-symmetry metallic phases as well as superconductors whose pairing mechanism and order parameter symmetry remain unsettled. Strikingly, experiments have revealed prominent new superconducting regions in rhombohedral bilayer and trilayer graphene devices with proximity-induced Ising spin-orbit coupling. We propose that these superconductors descend from a common spin-canted normal state that spontaneously breaks a U(1) spin symmetry and thus supports a gapless (Goldstone) magnon mode. In particular, we develop a scenario wherein pairing in the spin-canted state emerges from a novel type of magnon-mediated attraction: Contrary to conventional mechanisms that involve the exchange of a single boson, we show that second-order processes that exchange two magnons are dominant and produce an $s$-wave pairing interaction featuring a unique logarithmic low-frequency divergence. This low-frequency divergence disappears when spin-orbit coupling vanishes, providing a promising explanation for spin-orbit-enabled pairing. Numerous other experimental observations—including nontrivial dependence of superconductivity on the spin-orbit coupling strength, in-plane magnetic fields, and Fermi surface structure—also naturally follow from our scenario.
@@ I Introduction
Rhombohedral graphene—composed of layers stacked in an ‘ABC’ pattern—provides an extremely clean, gate-tunable platform for emergent quantum phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As prominent examples, ‘pure’ bilayer and trilayer devices (i.e., without proximity effects from adjacent 2D materials) exhibit exceptionally rich phase diagrams. Indeed, electrically varying the density and perpendicular displacement field reveals a series of correlated metallic states as well as unconventional superconductors [4, 5, 6, 7]. Despite detailed distinctions such as fermiology, normal-state order, and response to magnetic fields, these superconducting states share a common unifying ingredient: Each one appears tethered to a narrow density window close to a normal-state phase transition, prompting the investigation of pairing scenarios based on critical fluctuations [12, 13, 14, 15] and other mechanisms [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
Placing rhombohedral bilayers and trilayers proximate to a transition metal dichalcogenide (TMD) effectively boosts graphene’s spin-orbit coupling by orders of magnitude [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], in turn enriching the phenomenology in a very interesting way. Namely, new superconducting phases emerge that display the following salient features [37, 38, 41, 40, 39, 42]: $(i)$ Rather than residing near a phase boundary, they occur over a broad density range within a normal state that hosts two large majority Fermi pockets coexisting with some number of small minority pockets. $(ii)$ Their optimal critical temperatures $T_{c}$ far exceed those arising in ‘pure’ devices. And $(iii)$ their susceptibility to depairing by in-plane magnetic fields varies nontrivially across the superconducting regions. Numerous scenarios for how spin-orbit coupling promotes Cooper pairing in this setting have been introduced [37, 45, 46, 26, 27, 47, 48, 29]. Subsequent experiments on bilayers [39] further showed that, as the spin-orbit-coupling imparted by the TMD increases, the strongest new superconductor populates a diminished area in the phase diagram—yet curiously exhibits an enhanced maximal $T_{c}$.
@@
Figure 1: Summary of the proposed pairing mechanism. (a) Illustration of spin-canting order. The spin polarization ${\boldsymbol{n}}_{K}$ (${\boldsymbol{n}}_{K^{\prime}}$) in each valley is represented by a red (blue) arrow on the Bloch sphere. Ising spin-orbit interaction defines an out-of-plane quantization axis that points in opposite directions in the two valleys. Spin-canted order arises when an in-plane magnetic moment is spontaneously generated, leading to a non-trivial polar angle $\theta_{0}$ of the magnetization vectors and an arbitrarily-chosen azimuthal angle. Gapless magnon (Goldstone) modes associated with the spontaneous symmetry breaking are shown by black arrows. (b) The gapless magnons can mediate pairing in each Fermi surface through a second-order process that scatters Cooper pairs to the other band and back. This interaction yields an $s$-wave order parameter and is divergent at low frequencies, providing crucial retardation effects that benefit superconductivity.
@@
Given that the spin-canted phase evolves smoothly from a simple Stoner ferromagnet upon turning on $\lambda_{I}$, one can then ask why the former would be more conducive to Cooper pairing than the latter. A first hint surfaces from examining their collective modes. Both phases spontaneously break a continuous symmetry—U(1) spin rotations for the spin-canted phase and SU(2) for the Stoner ferromagnet—and hence both support gapless Goldstone magnon excitations. A qualitative distinction nevertheless exists: Because of the preceding symmetry distinction, the Goldstone mode disperses linearly with momentum for the spin-canted phase but quadratically with momentum for the Stoner ferromagnet. Might then low-energy magnon modes provide a pairing glue whose effectiveness depends critically on their dispersion?
We show that linearly dispersing magnons in the spin-canted phase can efficiently mediate attraction via a novel mechanism that differs sharply from conventional boson-mediated pairing. In the conventional picture, superconductivity emerges from Cooper pairs scattering via the exchange of a single boson (most commonly a phonon). Due to kinematic constraints, however, exchange of single low-energy magnons is ineffective at mediating pairing in the spin-canted phase. Nevertheless, we find that two-magnon exchange processes—wherein Cooper pairs near the Fermi surface scatter to virtual high-energy states and then back—gives rise to a pairing interaction that diverges at low frequency and favors $s$-wave superconductivity. This divergence encodes strong attraction over a narrow frequency range, implying significant retardation effects that are crucial for overcoming the pair-breaking tendencies of Coulomb repulsion. Interestingly, strong pairing arises here because two-magnon processes provide a way around a well-known constraint on the coupling strength between Goldstone modes and electrons [52] (see discussion in Sec. III.1). We additionally establish that two-magnon exchange events dominate over all higher-magnon processes, and that self-energy corrections do not spoil the pairing instability. Crucially, in the absence of spin orbital coupling, quadratically dispersing magnons characteristic of the simple SU(2) Stoner ferromagnet yield only a non-divergent pairing interaction. Two-magnon exchange operating within a spin-canted phase therefore naturally accounts for spin-orbit-enabled superconductivity in rhombohedral graphene.
Our proposed scenario also accommodates more refined experimerimental observations enumerated earlier. First, recall that spin-orbit-enabled superconductivity entails small minority Fermi pockets; moreover, in trilayers superconductivity abruptly turns off when they are vacated through a Lifshitz transition [41]. These observations hint that pairing may originate primarily from minority carriers. We predict that, due to the divergent nature of the magnon-mediated pairing interaction, its strength within a given Fermi surface scales inversely with the corresponding Fermi momentum—hence pairing indeed preferentially arises in small pockets.
Second, Ref. 39 observed a tunable and gradual increase of the maximal $T_{c}$ with induced Ising spin-orbit-coupling $\lambda_{I}$, coinciding with a counter-intuitive reduction in the phase space occupied by superconductivity. Our scenario recovers these seemingly at odds trends as follows. The gradual rise in achievable $T_{c}$ simply reflects the fact that spin-orbit coupling is crucial for producing a divergent, narrow-band pairing interaction as already noted. We find that the pairing interaction depends sensitively on the canting angle $\theta_{0}$ and scales roughly linearly with $\lambda_{I}$; see Fig. 2. Ising spin-orbit interaction simultaneously competes, however, against Hund’s coupling that is responsible for in-plane ferromagnetism and the associated Goldstone mode [3, 5]. Consequently, increasing $\lambda_{I}$ chips away at the phase space occupied by the spin-canting phase—as seen in Hartree-Fock simulations [50, 51]—and in doing so shrinks the parameter regime in which the magnon-mediated pairing glue thrives.
@@
Our theory may also explain unusual trends seen in the effects of in-plane magnetic fields on spin-orbit-enabled superconductivity. While naively one might anticipate high tolerance to in-plane magnetic fields due to Ising spin-orbit interaction [53, 54], experiments have measured critical in-plane fields that vary from well above the Pauli limit to below the Pauli limit depending on parameters [37, 38, 40, 41, 39]. Orbital or Rashba spin-orbit effects may at least partially explain these trends, though our scenario predicts a new depairing mechanism that could dominate the in-plane field response. Specifically, in-plane magnetic fields explicitly break the U(1) spin symmetry that the spin-canting phase breaks spontaneously—thereby gapping out the Goldstone mode, cutting off the low-frequency divergence in the pairing interaction, and sharply reducing the effectiveness of magnon-mediated pairing. This behavior stands in stark contrast to non-magnetic pairing glues, such as phonons, where superconductivity would be suppressed primarily by field-induced breaking of pairing resonances at the Fermi energy.
@@
Figure 2: Magnon modes and pairing interaction. The upper line illustrates the presumed phase diagram as a function of Ising spin-orbit coupling $\lambda_{I}$; a spin-canted phase appears up to a critical value $\lambda_{I}^{c}$, beyond which a symmetric spin-valley-locked phase emerges. The red curve sketches the magnon-mediated pairing interaction strength [see Eq. (50)], which depends sensitively on the canting angle $\theta_{0}$ and hence on spin-orbit coupling. Spheres illustrate the two relevant magnon modes in each regime: a valley-symmetric mode ($s-$, solid black arrow) and a valley-antisymmetric mode ($a-$, dashed gray arrow)—in valleys $K$ (blue arrows) and $K^{\prime}$ (red arrows). Insets (i)-(iv) show the associated magnon spectra. (i) In the symmetric spin-valley-locked phase, both magnon modes are gapped, and the pairing interaction is correspondingly suppressed. (ii) At the spin-canting phase transition both modes become gapless and have identical low-energy dispersion, and both contribute non-negligibly to the pairing interaction. (iii) In the canted phase where $\theta_{0}\neq 0$, the only gapless mode is the Goldstone mode $s-$ associated with the spontaneous symmetry breaking. (iv) When $\theta_{0}\rightarrow\pi/2$ (corresponding to vanishing spin-orbit interaction), the $s-$ mode acquires quadratic dispersion characteristic of an SU(2) ferromagnet. In this limit we show that the magnon-mediated pairing interaction is qualitatively suppressed.
@@
We organize the rest of this article as follows. Section II describes the phenomenology of the spin-canted phase in rhombohedral graphene, using both an effective Landau-theory description as well as a microscopic model from which we extract the magnon spectrum. Section III derives the form of magnon-mediated interactions between electrons, and Sec. IV then investigates the Cooper-pairing problem mediated by those interactions. Section V addresses the survival of pairing when incorporating self-energy corrections. Finally, Sec. VI compares our predictions to prior experimental observations and proposes several future directions.
@@ II.4 Magnon fluctuations
where $\delta{\boldsymbol{m}}_{\xi}(q)$ are magnon fluctuation fields. To a good approximation, the soft magnon modes of interest correspond to long-wavelength fluctuations of ${\boldsymbol{m}}_{\xi}$ that are orthogonal to the saddle-point value of $\bar{{\boldsymbol{m}}}_{\xi}$. Technically, fluctuations along $\bar{{\boldsymbol{m}}}_{\xi}$ and transverse to $\bar{{\boldsymbol{m}}}_{\xi}$ mix due the low symmetry of the problem, but the former can in principle be safely integrated out, yielding only quantitative modifications to the low-energy magnon spectrum. For simplicity we barbarically project onto the transverse fluctuations and write
@@
$-\sum_{\xi}\delta{\boldsymbol{m}}_{\xi}(q)\cdot[{\boldsymbol{s}}_{\xi}(-q)-\bar{{\boldsymbol{s}}}_{\xi}]\bigg{\}}$ (18)
@@
$-2\sum_{\xi}\chi_{2}(q)\delta m^{\theta}_{\xi}(q)\delta m^{\phi}_{\xi}(-q)\bigg{\}}.$ (19)
@@
In Appendix C we extract the low-lying magnon spectrum from poles of the magnon Green’s functions, i.e., frequencies for which $G_{b\eta}^{-1}(q)=0$. For generic canting angles satisfying $20<\theta_{0}<\pi/$ 2, only the $s-$ mode is gapless, with small-${{\boldsymbol{q}}}$ dispersion
@@
$c_{s}$ $\approx\sqrt{8DJ}\bar{m}\cos\theta_{0},$ (29)
valid in the physically motivated limit $J\ll V$ and for momenta below $q_{\rm max}\sim\frac{1}{V}\sqrt{\frac{J}{D}}\cos\theta_{0}$. Notice that the linear-dispersion regime shrinks to zero when approaching the SU(2) limit ($\theta_{0}\rightarrow\pi/2$), as expected. Estimating the quantity $VD$ through dimensional analysis as $\sim 1/k_{F}^{2}$, one finds $q_{\rm max}\sim\sqrt{\frac{J}{V}}\cos\theta_{0}k_{F}$. The momentum cutoff for linear dispersion is thus parametrically smaller than the Fermi momentum at $J\ll V$.
The structure of the $U_{s}(\nu)$ matrix moreover reveals that, at low frequencies ($\nu\lesssim\frac{\tilde{J}}{\gamma}\cos^{2}\theta_{0}$), $s-$ primarily encodes valley-symmetric $\delta m^{\phi}$ fluctuations, with subleading weight on $\delta m^{\theta}$ fluctuations. Physically, this Goldstone mode reflects long-wavelength azimuthal fluctuations in the orientation of the spontaneously chosen in-plane moment. One can therefore approximately project onto the sole low-energy magnon mode in this regime by simply replacing $\delta m^{\phi}_{\xi}\rightarrow\delta m_{s-}$ and $\delta m^{\theta}_{\xi}\rightarrow 0$.
At $\theta_{0}=0$, the $a-$ mode becomes gapless and degenerate with the $s-$ mode. This additional soft mode primarily involves (for $\nu\lesssim\tilde{J}/\gamma$) valley-antisymmetric $\delta m^{\theta}$ fluctuations. Note that $\theta_{0}=0$ corresponds to the phase boundary between canted and uncanted order, and hence no symmetries are broken in this limit. The $s-$ and $a-$ modes are not Goldstone modes, but rather respectively encode critical long-wavelength fluctuations in the in-plane moment along the $y$ and $x$ directions—reflecting proximity to spin-canted order.

### v3

@@ Abstract
Rhombohedral graphene multilayers host various broken-symmetry metallic phases as well as superconductors whose pairing mechanism and order parameter symmetry remain unsettled. Strikingly, experiments have revealed prominent new superconducting regions in rhombohedral bilayer and trilayer graphene devices with proximity-induced Ising spin-orbit coupling. We propose that these superconductors descend from a common spin-canted normal state that spontaneously breaks a U(1) spin symmetry and thus supports soft magnon modes. In particular, we show that these soft modes can mediate pairing through inter-band scattering events that are symmetry-forbidden in the absence of spin-orbit coupling, thus providing a promising explanation for spin-orbit-enabled pairing. Numerous other experimental observations—including nontrivial dependence of superconductivity on the spin-orbit coupling strength, in-plane magnetic fields, and Fermi surface structure—also naturally follow from our scenario.
@@ I Introduction
Rhombohedral graphene—composed of layers stacked in an ‘ABC’ pattern—provides an extremely clean, gate-tunable platform for emergent quantum phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. As prominent examples, ‘pure’ bilayer and trilayer devices (i.e., without proximity effects from adjacent 2D materials) exhibit exceptionally rich phase diagrams. Indeed, electrically varying the density and perpendicular displacement field reveals a series of correlated metallic states as well as unconventional superconductors [4, 5, 6, 7]. Despite detailed distinctions such as fermiology, normal-state order, and response to magnetic fields, these superconducting states share a common unifying ingredient: Each one appears tethered to a narrow density window close to a normal-state phase transition, prompting the investigation of pairing scenarios based on critical fluctuations [12, 13, 14, 15] and other mechanisms [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
Placing rhombohedral bilayers and trilayers proximate to a transition metal dichalcogenide (TMD) effectively boosts graphene’s spin-orbit coupling by orders of magnitude [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], in turn enriching the phenomenology in a very interesting way. Namely, new superconducting phases emerge that display the following salient features [40, 41, 44, 43, 42, 45]: $(i)$ Rather than residing near a phase boundary, they occur over a broad density range within a normal state that hosts two large majority Fermi pockets coexisting with some number of small minority pockets. $(ii)$ Their optimal critical temperatures $T_{c}$ far exceed those arising in ‘pure’ devices. And $(iii)$ their susceptibility to depairing by in-plane magnetic fields varies nontrivially across the superconducting regions. Numerous scenarios for how spin-orbit coupling promotes Cooper pairing in this setting have been introduced [48, 49, 28, 29, 50, 51, 30]. Recent experiments on bilayers [42] further showed that, as the spin-orbit coupling imparted by the TMD increases, the strongest new superconductor populates a diminished area in the phase diagram—yet curiously exhibits an enhanced maximal $T_{c}$.
@@
Figure 1: Summary of the proposed pairing mechanism. (a) Illustration of spin-canting order. The spin polarization ${\boldsymbol{n}}_{K}$ (${\boldsymbol{n}}_{K^{\prime}}$) in each valley is represented by a red (blue) arrow on the Bloch sphere. Ising spin-orbit interaction defines an out-of-plane quantization axis that points in opposite directions in the two valleys. Spin-canted order arises when an in-plane magnetic moment is spontaneously generated, leading to a non-trivial polar angle $\theta_{0}$ of the magnetization vectors and an arbitrarily chosen azimuthal angle. Gapless magnon (Goldstone) modes associated with the spontaneous symmetry breaking are shown by black arrows. Soft magnon modes can mediate pairing through interband (spin-flipping) scattering, depicted with (b) Feynman diagrams and (c) kinematics.
@@
Given that the spin-canted phase evolves smoothly from a simple Stoner ferromagnet upon turning on $\lambda_{I}$, one can then ask why the former would be more conducive to Cooper pairing than the latter. A first hint surfaces from examining their collective modes. Both phases spontaneously break a continuous symmetry—U(1) spin rotations for the spin-canted phase and SU(2) for the Stoner ferromagnet—and hence both support soft order parameter fluctuations: a gapless Goldstone magnon excitation, and a spin-canting mode with parametrically small gap that becomes critical at the canting transition. A qualitative distinction nevertheless exists. Namely, the simple Stoner ferromagnet preserves a U(1) subgroup of SU(2) spin rotations, and hence the magnetization must be preserved by arbitrary magnon-mediated processes. Might then soft magnon modes provide a pairing glue whose effectiveness depends critically on the symmetry-allowed scattering events that they mediate?
In the case with spin-orbit coupling, we indeed show that magnons that are softened near or inside the spin-canted phase can mediate attraction through spin-flipping, interband scattering that favors $s$-wave pairing with an opposite sign between the majority and minority Fermi surfaces [^1] Throughout this paper, we refer to a “pseudospin-singlet” $s$-wave channel where the pairing gap $\Delta({\bf p})$ has no sign alternation when ${\bf p}$ moves along each Fermi surface. However, due to fermion anticommutation relation, $\Delta({\bf p})$ changes sign when $\bf p$ moves between valleys. When viewed in the whole hexagonal Brillouin zone, this pairing channel can thus be classified as $f$-wave.. The energy of the soft magnon modes (over the relevant momentum range for inter-band scattering) constitutes a natural low-frequency scale, providing retardation effects that are crucial to overcome the pair-breaking tendencies of Coulomb repulsion. In the absence of spin-orbit coupling, by contrast, magnons of the SU(2) Stoner ferromagnet yield a vanishing inter-band pairing interaction due to constraints imposed by the remaining U(1) symmetry, thus furnishing a plausible explanation for spin-orbit-enabled superconductivity in rhombohedral graphene.
Our scenario also accommodates more refined experimental observations enumerated earlier. First, recall that spin-orbit-enabled superconductivity entails minority Fermi pockets; moreover, in trilayers superconductivity abruptly turns off when they are vacated through a Lifshitz transition [44]. These observations hint that pairing requires the presence of minority carriers, in harmony with spin-flipping scattering that acts between the majority and minority Fermi surfaces. Second, Ref. 42 observed a tunable and gradual increase of the maximal $T_{c}$ with induced Ising spin-orbit-coupling $\lambda_{I}$, coinciding with a reduction in the phase space occupied by superconductivity. The concurrence of these two trends is surprising because in a conventional scenario, where the pairing glue is non-magnetic (e.g., phonon-mediated), spin-orbit coupling is expected to either uniformly help or hurt superconductivity through its effect on the Fermi surface. Our scenario, however, naturally recovers these seemingly at odds trends: Increasing $\lambda_{I}$ decreases the canting angle $\theta_{0}$ [see Eq. (4)]—in turn enhancing the magnon-mediated pairing interaction [see Eq. (38)] and thereby boosting $T_{c}$. At the same time, spin-orbit interaction competes against Hund’s coupling, such that increasing $\lambda_{I}$ erodes the phase space occupied by the spin-canted phase, whose Goldstone modes provide the pairing glue in our scenario. Increasing $\lambda_{I}$ thus also narrows the superconducting phase. Recent Hartree-Fock simulations support the predicted change in $\theta_{0}$ with $\lambda_{I}$ [54] accounting for the rise in $T_{c}$, while also finding that the phase space for spin-canting order evolves with $\lambda_{I}$ similarly to the phase space for superconductivity in experiment (see Fig. 7 in Ref. 53 and Fig. 4 in Ref. 54). To the best of our knowledge, no other pairing mechanism relate superconductivity to spin canting in these systems.
@@
Our theory may also explain unusual trends seen in the effects of in-plane magnetic fields on spin-orbit-enabled superconductivity. While naively one might anticipate high tolerance to in-plane magnetic fields due to Ising spin-orbit interaction [55, 56], experiments have measured critical in-plane fields that vary from well above the Pauli limit to below the Pauli limit depending on parameters [40, 41, 43, 44, 42]. Orbital or Rashba spin-orbit effects may at least partially explain these trends, though our scenario predicts a new depairing mechanism that could contribute to the in-plane field response. Specifically, in-plane magnetic fields explicitly break the U(1) spin symmetry that the spin-canted phase breaks spontaneously—thereby gapping out the Goldstone mode and reducing the effectiveness of magnon-mediated pairing. This behavior stands in contrast to non-magnetic pairing glues, such as phonons, where superconductivity would be suppressed primarily by field-induced breaking of pairing resonances at the Fermi energy.
@@
Figure 2: Magnon modes and pairing interaction. The upper line illustrates the presumed phase diagram as a function of Ising spin-orbit coupling $\lambda_{I}$; a spin-canted phase appears up to a critical value $\lambda_{I}^{c}$, beyond which a symmetric spin-valley-locked phase emerges. The red curve sketches the magnon-mediated pairing interaction strength [see Eq. (45)], which depends sensitively on the canting angle $\theta_{0}$ and hence on spin-orbit coupling. Spheres illustrate the two relevant low-energy magnon modes in each regime: a valley-symmetric mode ($s-$, solid black arrow) and a valley-antisymmetric mode ($a-$, dashed gray arrow)—in valleys $K$ (blue arrows) and $K^{\prime}$ (red arrows). Insets (i)-(iv) show the associated magnon spectra. (i) In the symmetric spin-valley-locked phase, both magnon modes are gapped, and the pairing interaction is correspondingly suppressed. (ii) At the spin-canted phase transition both modes become gapless and have identical low-energy dispersion.(iii) In the canted phase where $\theta_{0}\neq 0$, the only gapless mode is the Goldstone mode $s-$ associated with the spontaneous symmetry breaking; however, the $a-$ magnon gap is parametrically small in $J/V$ and also contributes appreciably to pairing. (iv) When $\theta_{0}\rightarrow\pi/2$ (corresponding to vanishing spin-orbit interaction), the $s-$ mode acquires quadratic dispersion characteristic of an SU(2) ferromagnet. In this limit we show that magnon-mediated interband pairing vanishes by symmetry.
@@
We organize the rest of this article as follows. Section II describes the phenomenology of the spin-canted phase in rhombohedral graphene, using both an effective Landau-theory description as well as a microscopic model from which we extract the magnon spectrum. Section III derives the form of magnon-mediated interactions between electrons, and Sec. IV then investigates the Cooper-pairing problem mediated by those interactions. Finally, Sec. V compares our results to prior experimental observations, makes key testable predictions, and proposes several future directions.
@@ II.4 Magnon fluctuations
where $\delta{\boldsymbol{m}}_{\xi}(q)$ are magnon fluctuation fields. To a good approximation, the soft magnon modes of interest correspond to long-wavelength fluctuations of ${\boldsymbol{m}}_{\xi}$ that are orthogonal to the saddle-point value of $\bar{{\boldsymbol{m}}}_{\xi}$. Technically, fluctuations along $\bar{{\boldsymbol{m}}}_{\xi}$ and transverse to $\bar{{\boldsymbol{m}}}_{\xi}$ mix due the low symmetry of the problem, but the former can in principle be safely integrated out, yielding only quantitative modifications to the low-energy magnon spectrum. For simplicity we barbarically project onto the transverse fluctuations [^2] While this procedure is safe up to leading (linear) order in the fermion-magnon coupling, it throws out corrections to quadratic order in $\delta m$ that become important when considering higher-order magnon exchange processes. In fact, these corrections are crucial in ensuring that the intra-band electron-Goldstone mode coupling vanishes when ${\boldsymbol{q}}\rightarrow 0$, as expected from general considerations [58]. and write
@@
$-\sum_{\xi}\delta{\boldsymbol{m}}_{\xi}(q)\cdot[{\boldsymbol{s}}_{\xi}(-q)-\bar{{\boldsymbol{s}}}_{\xi}]$ (18)
@@
$-2\sum_{\xi}\chi_{2}(q)\delta m^{\theta}_{\xi}(q)\delta m^{\phi}_{\xi}(-q).$ (19)
@@
In Appendix C we extract the low-lying magnon spectrum from poles of the magnon Green’s functions, i.e., frequencies for which $G_{b\eta}^{-1}(q)=0$. As detailed there, the $a+$ and $s+$ modes exhibit gaps comparable to the Fermi energy $\epsilon_{F}$. We therefore focus below on the other two modes, $s-$ and $a-$, that are much softer (see also Fig. 2). For generic canting angles satisfying $20<\theta_{0}<\pi/$ 2, only the $s-$ mode is gapless. The structure of the $U_{s}(\nu)$ matrix reveals that, at low frequencies ($\nu\lesssim\frac{\tilde{J}}{\gamma}\cos^{2}\theta_{0}$), $s-$ primarily encodes valley-symmetric $\delta m^{\phi}$ fluctuations, with subleading weight on $\delta m^{\theta}$ fluctuations. Physically, this Goldstone mode reflects long-wavelength azimuthal fluctuations in the orientation of the spontaneously chosen in-plane moment. The small-${{\boldsymbol{q}}}$ dispersion reads
@@
$c_{s}$ $=\sqrt{\frac{D}{\kappa z_{s}}}\cos\theta_{0},$ (29)
where we defined the dimensionless quantity $z_{s}=\cos^{2}\theta_{0}+\frac{V-J}{2J}$. The linear-dispersion regime is valid for momenta below $q_{\rm max}\sim\frac{1}{V}\sqrt{\frac{J}{D}}\cos\theta_{0}$ and thus shrinks to zero when approaching the SU(2) limit ($\theta_{0}\rightarrow\pi/2$), as expected. Estimating the quantity $VD$ through dimensional analysis as $\sim 1/k_{F}^{2}$, one finds $q_{\rm max}\sim\sqrt{\frac{J}{V}}\cos\theta_{0}k_{F}$. Given that we expect $V/J\sim 10$, the momentum cutoff for linear dispersion is comparable to the Fermi momentum. See Appendix C for further numerical investigation of the magnon dispersion.
The dispersion of the $a-$ mode, which at low frequencies primarily involves valley-antisymmetric $\delta m^{\theta}$ fluctuations, is given by
$\omega_{a-}({{\boldsymbol{q}}})$ $=\sqrt{\Omega_{a-}^{2}+c_{a}^{2}{{\boldsymbol{q}}}^{2}},$ (30)
$c_{a}$ $=\sqrt{\frac{D}{\kappa z_{a}}}(1+\sin^{2}\theta_{0})^{1/2},$ (31)
where we defined $z_{a}=(1+\sin^{2}\theta_{0})+\frac{V-J}{2J}$ and the magnon gap $\Omega_{a-}=\sqrt{\frac{2\tilde{J}}{z_{a}\kappa}}\sin\theta_{0}$. Up to leading order in $J/V$, we obtain that $\Omega_{a-}\sim\left(\frac{J\sin\theta_{0}}{V}\right)\sqrt{\frac{1}{V\kappa}}\sim\left(\frac{J\sin\theta_{0}}{V}\right)\bar{m}$. When the polarization densities are comparable to the total electronic density [40, 41, 44, 43, 42, 45], we further have $\bar{m}\sim\epsilon_{F}$. Consequently, the $a-$ magnon gap lies well below $\epsilon_{F}$ in the experimentally relevant regime as claimed.
At $\theta_{0}=0$, the $a-$ mode becomes gapless and degenerate with the $s-$ mode. Note that $\theta_{0}=0$ corresponds to the phase boundary between canted and uncanted order, and hence no symmetries are broken in this limit. The $s-$ and $a-$ modes are not Goldstone modes, but rather respectively encode critical long-wavelength fluctuations in the in-plane moment along the $y$ and $x$ directions—reflecting proximity to spin-canted order.